# Tree-tubings and the combinatorics of resurgent Dyson–Schwinger equations

### v1

@@ Acknowledgements
KY thanks Nick Olson-Harris, and GD thanks Ovidiu Costin, for useful discussions.
@@ 2. Solving Dyson-Schwinger equations by tubings
The number of tubes in a tubing $\tau$ of a tree $t$ with $n$ vertices is $12n-$ 1. This can be proved by a quick induction on the binary structure of the tubing. Another useful thing to note is that for a fixed tubing $\tau$ every edge of the graph corresponds to a parititioning of some tube into two subtubes in $\tau$, and every such partitioning in $\tau$ corresponds to an edge of the tree.
We assume, in accordance to the physical origins of (1), that the Mellin transform $F(\rho)$ has a simple pole at the origin and that we can write it in terms of its Laurent series expansion
@@ 3. Combinatorial interpretation of the anomalous dimension
where we used that $\sum_{v\in V(t)}b(v,\tau)$ is the number of tubes of $\tau$, namely $12|t|-$ 1, since every tube has a unique root and $N(t)$ is the number of tubings of $t$.
After pulling out one factor of $2$ and scaling the coupling $x$ by $1/4$, this means that the anomalous dimension becomes simply the generating series for tubings. By [1] Section 5 there is a size preserving bijection between tubings of plane rooted trees and rooted connected chord diagrams. Therefore, after scaling, the anomalous dimension in the Yukawa case is the generating series for rooted connected chord diagrams. This was observed without proof in [7], and was an important insight in [5], as this combinatorial interpretation for the anomalous dimension allowed for a resummation of all the instantons. For this resummation it was critical that the generating function of rooted connected chord diagrams fulfills a simple functional equation. A resurgence-inspired framework put forward by the first author based on generating functions and a differential operator that extracts asymptotic information from such functional relations [3] was critical to achieve this. This framework led to complete asymptotic information on the functional equation for the generating function of rooted connected chord diagrams and thereby to a resummable all-instanton solution.
@@
where $a_{1},\ldots,a_{m}\in\mathbb{Z}$ and $0<a_{1}\leq a_{2}\leq\cdots\leq a_{m}$. For the rest of the paper we will assume our Mellin transform is of this form. To return to a more physical scaling and thus agree with past work, see Remark 3.11. We know from elementary enumeration that the generating series for $k$-ary strings (that is, strings or words from an alphabet of $k$ letters) is $1/(1-kx)$. We also know from elementary enumeration that if $A(x)$ and $B(x)$ are generating series for $\mathcal{A}$ and $\mathcal{B}$ respectively then $A(x)B(x)$ is the generating series for $\mathcal{A}\times\mathcal{B}$ where the size of an ordered pairs is the sum of the sizes of its two components [13, Part. A]. Additionally, suppose we have an ordered pair of a word on the alphabet $\Omega$ and a word on the alphabet $\Omega^{\prime}$ where $\Omega\cap\Omega^{\prime}=\emptyset$. By concatenating the two words, we can view any such ordered pair as a single word on the alphabet $\Omega\cup\Omega^{\prime}$ where all letters from $\Omega$ must appear before all letters from $\Omega^{\prime}$. Taking all this together we get that (5), after multiplying by $\rho$ to remove the pole, is the generating series (in $\rho$) for words with letters from $\Omega_{1}\cup\Omega_{2}\cup\cdots\Omega_{m}$ where the $\Omega_{i}$ are disjoint alphabets, $\Omega_{i}$ has $a_{i}$ letters, and where the words have all $\Omega_{i}$ letters before any $\Omega_{j}$ letters for all $i<j$. Given such $\Omega_{i}$, call a word on $\Omega_{1}\cup\cdots\cup\Omega_{m}$ where all letters from $\Omega_{i}$ occur before any letters from $\Omega_{j}$ for each $i<j$ an admissible word.
@@
Now label each upper tube with a letter from $\Omega_{1}\cup\cdots\cup\Omega_{m}$ with the restriction that whenever $i<j$ and we have a letter from $\Omega_{i}$ and a letter from $\Omega_{j}$ which label two tubes with the same root, then the letter from $\Omega_{i}$ is on a tube which is inside the tube with the letter from $\Omega_{j}$. With these constraints, the letters on tubes with the same root give an admissible word when the letters are formed into a word by the innermost to outermost order of the tubes, and every way of assigning admissible words of length $b(v,\tau)$ to each vertex is obtained in this way. [^1] We could just as well note that every vertex is in one tube of size $1$ in $\tau$ and so also $b(v,\tau)-1$ is the number of tubes of size $>1$ of $\tau$ for which $v$ is the root. Thus we could replace “upper tube” with “tube of size $>1$” in the labelling by letters and obtain all the same results. However, as shown in [18], the upper tubes are structurally nicer, and so are a preferable choice.
Encapsulate such a labelling of the tubes of a tubing in a definition.
@@
Let $\Omega_{1},\ldots\Omega_{m}$ be disjoint alphabets with $|\Omega_{i}|=a_{i}$. Let $\tau$ be a tubing of a plane rooted tree $t$. Call a labelling of the upper tubes of $\tau$ with letters from $\Omega_{1}\cup\cdots\cup\Omega_{m}$a compatible labelling (or a $a_{1},\ldots,a_{m}$-compatible labelling, or a $\Omega_{1},\ldots,\Omega_{m}$-compatible labelling, when distinguishing the sizes of alphabets is necessary) if whenever $i<j$ and there is a letter from $\Omega_{i}$ and a letter from $\Omega_{j}$ which label two tubes with the same root, then the letter from $\Omega_{i}$ is on a tube which is inside the tube with the letter from $\Omega_{j}$.
Let $L(\tau)$ be the number of compatible labellings of $\tau$. Then continuing the calculation of $\gamma(x)$
@@
In particular, $\gamma(x)$ is the generating series for compatible labellings of tubings of plane rooted trees.
@@
Let $\Omega_{1},\ldots\Omega_{m}$ be disjoint alphabets with $|\Omega_{i}|=a_{i}$. The anomalous dimension of the perturbative solution to (1) with Mellin transform (5) is the generating series for compatible labellings of tubings of plane rooted trees.
@@
$\frac{1}{\rho(3-\rho)(2-\rho)(1-\rho)}=\frac{1}{6\rho(1-\frac{1}{3}\rho)(1-\frac{1}{2}\rho)(1-\rho)}=\frac{1}{36\rho^{\prime}(1-2\rho^{\prime})(1-3\rho^{\prime})(1-6\rho^{\prime})}$
@@
Proposition 3.3 then tells us that, after these scalings, [^2] See Remark 3.11 on reversing the scalings in order to compare with [8] and [6] directly. the anomalous dimension for the perturbative solution to the Dyson-Schwinger equation is the generating series of compatible labellings of tubings of plane rooted trees with three alphabets, one with 2 letters, one with 3 letters and one with 6 letters with the compatibility restriction, so tubes with letters from smaller alphabets appear inside tubes with the same root and with letters from larger alphabets.
In particular, $2,3,6$-compatible labellings of tubings of plane rooted trees give a combinatorial interpretation for the sequence given at the end of section 2 of [6] which the authors observed had no known interpretation at the time.
To make the tubing construction more concrete let us work out the first few terms explicitly. The tree with a single vertex has a single tubing consisting of just one tube. This tube is by convention a lower tube and so does not get a letter. Therefore the coefficient of $x^{1}$ in $\gamma(x)$ is $1$ in this model. There is one tree with two vertices and one tubing of it, illustrated on the left in Figure 1. The only upper tube in this case is the tube containing only the root, so there is one upper tube rooted at the root and no upper tubes rooted at the other vertex. These numbers are illustrated in the figure. Thus there is only one tube that gets a letter, and it can have any of the $2+3+6=11$ letters, so the coefficient of $x^{2}$ in $\gamma(x)$ is $11$.
@@
Next consider the coefficient of $x^{3}$ in $\gamma(x)$. There are two rooted trees on three vertices, each of which has two tubings as illustrated on the right in Figure 1. The number of upper tubes rooted at each vertex is indicated in the figure. For the third tubing from the right, there are $11$ ways to label each upper tube since they are rooted at different vertices so there is no constraint on which letters can be chosen. For each of the other three tubings we must label a pair of tubes with the same root. If the outer of the two tubes is labelled by a letter from $\Omega_{3}$ then there is no constraint on the label of the inner of the two tubes. If the outer of the two tubes is labelled by a letter from $\Omega_{2}$ then the label of the inner of the two tubes must be from $\Omega_{2}\cup\Omega_{3}$ while if the label of the outer of the two tubes is from $\Omega_{1}$ then the label of the outer of the two tubes must also be from $\Omega_{1}$. Consequently there are $6\cdot 11+3\cdot(2+3)+2\cdot 2=85$ compatible labellngs of such a tubing.
All together this gives that the coefficient of $x^{3}$ in $\gamma(x)$ is $37611^{2}+3\cdot 85=$ 376, agreeing with the explicit calculations of [8, 6].
@@
To build this system of differential equations, first consider how to recursively build tubings. As discussed in [1], by the definition of tubing, the outermost tube of any tubing is partitioned into subtubes and each of those subtubes along with all the tubes within determine a tubing of a connected subgraph of the original tree. Furthermore, take any two tubings $\tau_{1}$ and $\tau_{2}$ of plane rooted trees $t_{1}$ and $t_{2}$ respectively and then choose an insertion place in $t_{2}$. Adding an edge at the insertion point connecting to the root of $t_{1}$ builds a new tree $t$ and taking all the tubes of $\tau_{1}$ and $\tau_{2}$ along with the outer tube for $t$ gives a tubing $\tau$ of $t$, and every tubing of every rooted tree can be formed uniquely in this way. Furthermore, if a plane rooted tree has $n$ vertices then it had $12n-$ 1 insertion places. Therefore, if $N(n)$ is the number of tubings of rooted trees on $n$ vertices then
@@
The differential operator $12x\frac{d}{dx}-$ 1 serves to pull out the coefficient $12i-$ 1, and hence serves to count the insertion places. This is a special case of the general interpretation of $x\frac{d}{dx}$ as a pointing operator in enumerative combinatorics [13, Ch. A.I]. The recurrence (6) for tubings of plane rooted trees was used in [1] to give a bijection between tubings of plane rooted trees and rooted connected chord diagrams, which are classically known to have the same recurrence [16].
We will use this same construction but keeping track of compatible labellings to prove the following theorem.
@@
Let $\Omega_{1},\ldots\Omega_{m}$ be disjoint alphabets with $|\Omega_{i}|=a_{i}$. Let $L_{j}(x)$ be the generating series for $\Omega_{1},\ldots\Omega_{m}$-compatible labellings of tubings of plane rooted trees for which the label on the outermost tube comes from $\Omega_{j}$. Then the anomalous dimension $\gamma(x)$ of the perturbative solution to (1) with Mellin transform (5) satisfies the following system of differential equations.
@@
As before let $\tau_{1}$ and $\tau_{2}$ be compatibly labelled tubings of plane rooted trees $t_{1}$ and $t_{2}$ respectively and choose an insertion place in $t_{2}$. Construct $\tau$ and $t$ as above by inserting $t_{1}$ into $t_{2}$ at the chosen insertion place. The labellings of the tubes of $\tau_{1}$ and $\tau_{2}$ remains compatible in $\tau$. The only thing missing is a label for the upper tube of the partition of the outermost tube, which had previously been a lower tube and is now an upper tube. Call this label the outermost label of $\tau$. The outermost label of $\tau$ is constrained to be from an alphabet $\Omega_{j}$ with $j\geq i$ where the outermost label of $\tau_{2}$ is from $\Omega_{i}$, or if $t_{2}$ is a single vertex and so has no labelled tubes then the outermost label of $\tau$ is unconstrained. Furthermore, this is the only constraint on the outermost label of $\tau$ and every compatible labelling can be obtained uniquely in this way. The labels from $\tau_{1}$ put no constraint on the outermost label of $\tau$ since they do not share roots.
As above we can encode this decomposition by generating series. Every compatible labelling of a tubing of a plane rooted tree has its outermost label from exactly one of the alphabets except for the tree with one vertex and so
@@
For a compatible labelling with the outermost label from $\Omega_{j}$, the decomposition described above tells us we have $a_{j}$ choices for the outermost label, we can take any compatible labelling of a tubing of a plane rooted tree for $\tau_{1}$, hence giving a factor of $\gamma(x)$ in the generating series, and we can take either a compatible labelling of a tree with one vertex or any compatible labelling with outermost label from $\Omega_{1}\cup\cdots\cup\Omega_{j}$ for $\tau_{2}$, where as before the differential operator $12x\frac{d}{dx}-$ 1 incorporates the count of the number of insertion places, and so
@@
Let $\gamma_{j}:=x+\sum_{k=1}^{j}L_{k}$. For $j=1$, we have $x=\gamma_{1}-L_{1}$, and since $L_{1}=a_{1}\gamma\left(2x\frac{d}{dx}-1\right)\gamma_{1}$ we have $x=\left[1-a_{1}\gamma\left(2x\frac{d}{dx}-1\right)\right]\gamma_{1}$. Similarly, $\gamma_{j+1}-L_{j+1}=\gamma_{j}$ implies $\left[1-a_{j+1}\gamma\left(2x\frac{d}{dx}-1\right)\right]\gamma_{j+1}=\gamma_{j}$. The result follows by induction, noting that $\gamma_{m}=\gamma$, by definition. ∎
@@
An advantage of the differential equation system form in Theorem 3.5 is that it provides a convenient way to compute the coefficients of the power series $\gamma(x)=\sum_{n\geq 1}g_{n}x^{n}$. To do so, we substitute power series ansätze for $\gamma(x)$ and $L_{j}(x)$ into the differential equation system. This yields a system of nonlinear difference equations for the coefficients that can be solved recursively (see Section 3.1.2 for a simple example of this procedure). We implemented this recursion as the python program gamma.py which is included in the ancillary files to this article. In a short time, this program computes the numbers $g_{n}$ for various input parameters $a_{1},\ldots,a_{m}\in{\mathbb{Z}}$ up to $n\approx 1000$.
@@
Its unique power series solution (see [17, A000699]) starts with
@@
This is one of the rare cases where the prefactor $\frac{1}{e}$ is known exactly thanks to a functional equation that the associated generating function fulfills (see [3, § 7.1]).
@@
This system is equivalent to a third order nonlinear ODE with quartic nonlinearity as in Corollary 3.6:
@@
The first coefficients of the unique power series solution (see [17, A051862]) are
@@
It is equivalent to a second order ODE with cubic nonlinearity,

### v3

@@ Acknowledgements
KY thanks Nick Olson-Harris, and GD thanks Ovidiu Costin, for useful discussions. We also thank Paul Balduf, Marc Bellon, David Broadhurst, and Dirk Kreimer for their helpful comments.
@@ 2. Solving Dyson-Schwinger equations by tubings
The number of tubes in a tubing $\tau$ of a tree $t$ with $n$ vertices is 2n12n-1. This can be proved by a quick induction on the binary structure of the tubing. Another useful thing to note is that for a fixed tubing $\tau$ every edge of the graph corresponds to a partitioning of some tube into two subtubes in $\tau$, and every such partitioning in $\tau$ corresponds to an edge of the tree.
The logarithmic divergence of the Mellin transformed Feynman diagram implies that $F(\rho)$ has a simple pole at the origin. So, in accordance to the physical origins of (1), we will assume that $F(\rho)$ can be written as the following Laurent series expansion
@@
Remark 2.1.
Symmetries of specific quantum field theories can further constrain the structure of the Laurent series $F(\rho)$. For instance, the conformal symmetry of Yukawa theory implies that $F(\rho)=F(2-\rho)$. Here, we only assume that $F(\rho)$ is of the general form above.
@@ 3. Combinatorial interpretation of the anomalous dimension
where we used that $\sum_{v\in V(t)}b(v,\tau)$ is the number of tubes of $\tau$, namely 2|t|12|t|-1, since every tube has a unique root and $N(t)$ is the number of tubings of $t$.
After pulling out one factor of $2$ and scaling the coupling $x$ by $1/4$, this means that the anomalous dimension becomes simply the generating series for tubings. By [3] Section 5 there is a size-preserving bijection between tubings of plane rooted trees and rooted connected chord diagrams. Therefore, after scaling, the anomalous dimension in the Yukawa case is the generating series for rooted connected chord diagrams. This was observed without proof in [10], and was an important insight in [8], as this combinatorial interpretation for the anomalous dimension allowed for a resummation of all the instantons. For this resummation it was critical that the generating function of rooted connected chord diagrams fulfills a simple functional equation. A resurgence-inspired framework put forward by the first author based on generating functions and a differential operator that extracts asymptotic information from such functional relations [6] was critical to achieve this. This framework led to complete asymptotic information on the functional equation for the generating function of rooted connected chord diagrams and thereby to a resummable all-instanton solution.
@@
where $a_{1},\ldots,a_{m}\in\mathbb{Z}$ and $0<a_{1}\leq a_{2}\leq\cdots\leq a_{m}$. For the rest of the paper we will assume our Mellin transform is of this form. To return to a more physical scaling and thus agree with past work, see Remark 3.12. We know from elementary enumeration that the generating series for $k$-ary strings (that is, strings or words from an alphabet of $k$ letters) is 1/(1kx)1/(1-kx). We also know from elementary enumeration that if $A(x)$ and $B(x)$ are generating series for $\mathcal{A}$ and $\mathcal{B}$ respectively then $A(x)B(x)$ is the generating series for $\mathcal{A}\times\mathcal{B}$ where the size of an ordered pairs is the sum of the sizes of its two components [17, Part. A]. Additionally, suppose we have an ordered pair of a word on the alphabet $\Omega$ and a word on the alphabet $\Omega^{\prime}$ where $\Omega\cap\Omega^{\prime}=\emptyset$. By concatenating the two words, we can view any such ordered pair as a single word on the alphabet $\Omega\cup\Omega^{\prime}$ where all letters from $\Omega$ must appear before all letters from $\Omega^{\prime}$. Taking all this together we get that (5), after multiplying by $\rho$ to remove the pole, is the generating series (in $\rho$) for words with letters from $\Omega_{1}\cup\Omega_{2}\cup\cdots\Omega_{m}$ where the $\Omega_{i}$ are disjoint alphabets, $\Omega_{i}$ has $a_{i}$ letters, and where the words have all $\Omega_{i}$ letters before any $\Omega_{j}$ letters for all $i<j$. Given such $\Omega_{i}$, call a word on $\Omega_{1}\cup\cdots\cup\Omega_{m}$ where all letters from $\Omega_{i}$ occur before any letters from $\Omega_{j}$ for each $i<j$ an admissible word.
@@
Note that we have a formula for $w_{n}$, but this will not be crucial in what follows. Namely, by the required order on the letters in an admissible word, we can sum over all possibilities for which $\Omega_{i}$ each letter is in by summing over $1\leq i_{1}\leq i_{2}\leq\cdots\leq i_{n}\leq m$, and for each position then we simply need to take the size of the alphabet, giving
(6) $w_{n}=\sum_{1\leq i_{1}\leq i_{2}\leq\cdots\leq i_{n}\leq m}a_{i_{1}}a_{i_{2}}\cdots a_{i_{n}}.$
@@
Now label each upper tube with a letter from $\Omega_{1}\cup\cdots\cup\Omega_{m}$ with the restriction that whenever $i<j$ and we have a letter from $\Omega_{i}$ and a letter from $\Omega_{j}$ which label two tubes with the same root, then the letter from $\Omega_{i}$ is on a tube which is inside the tube with the letter from $\Omega_{j}$. With these constraints, the letters on tubes with the same root give an admissible word when the letters are formed into a word by the innermost to outermost order of the tubes, and every way of assigning admissible words of length $b(v,\tau)$ to each vertex is obtained in this way. [^1] We could just as well note that every vertex is in one tube of size $1$ in $\tau$ and so also $b(v,\tau)-1$ is the number of tubes of size >1>1 of $\tau$ for which $v$ is the root. Thus we could replace “upper tube” with “tube of size >1>1” in the labeling by letters and obtain all the same results. However, as shown in [22], the upper tubes are structurally nicer, and so, they are a preferable choice.
Encapsulate such a labeling of the tubes of a tubing in a definition.
@@
Let $\Omega_{1},\ldots\Omega_{m}$ be disjoint alphabets with $|\Omega_{i}|=a_{i}$. Let $\tau$ be a tubing of a plane rooted tree $t$. Call a labeling of the upper tubes of $\tau$ with letters from $\Omega_{1}\cup\cdots\cup\Omega_{m}$a compatible labeling (or a $a_{1},\ldots,a_{m}$-compatible labeling, or a $\Omega_{1},\ldots,\Omega_{m}$-compatible labeling, when distinguishing the sizes of alphabets is necessary) if whenever $i<j$ and there is a letter from $\Omega_{i}$ and a letter from $\Omega_{j}$ which label two tubes with the same root, then the letter from $\Omega_{i}$ is on a tube which is inside the tube with the letter from $\Omega_{j}$.
Let $L(\tau)$ be the number of compatible labelings of $\tau$. Then continuing the calculation of $\gamma(x)$
@@
In particular, $\gamma(x)$ is the generating series for compatible labelings of tubings of plane rooted trees.
@@
Let $\Omega_{1},\ldots\Omega_{m}$ be disjoint alphabets with $|\Omega_{i}|=a_{i}$. The anomalous dimension of the perturbative solution to (1) with Mellin transform (5) is the generating series for compatible labelings of tubings of plane rooted trees.
@@
$\frac{1}{\rho(3-\rho)(2-\rho)(1-\rho)}=\frac{1}{6\rho(1-\frac{1}{3}\rho)(1-\frac{1}{2}\rho)(1-\rho)}$
$=\frac{1}{36\rho^{\prime}(1-2\rho^{\prime})(1-3\rho^{\prime})(1-6\rho^{\prime})}$
@@
Proposition 3.3 then tells us that, after these scalings, [^2] See Remark 3.12 on reversing the scalings in order to compare with [11] and [9] directly. the anomalous dimension for the perturbative solution to the Dyson-Schwinger equation is the generating series of compatible labelings of tubings of plane rooted trees with three alphabets, one with 2 letters, one with 3 letters and one with 6 letters with the compatibility restriction, so tubes with letters from smaller alphabets appear inside tubes with the same root and with letters from larger alphabets.
In particular, $2,3,6$-compatible labelings of tubings of plane rooted trees give a combinatorial interpretation for the sequence given at the end of section 2 of [9] which the authors observed had no known interpretation at the time.
To make the tubing construction more concrete let us work out the first few terms explicitly. The tree with a single vertex has a single tubing consisting of just one tube. This tube is by convention a lower tube and so, it does not get a letter. Therefore the coefficient of $x^{1}$ in $\gamma(x)$ is $1$ in this model. There is one tree with two vertices and one tubing of it, illustrated on the left in Figure 1. The only upper tube in this case is the tube containing only the root, so there is one upper tube rooted at the root and no upper tubes rooted at the other vertex. These numbers are illustrated in the figure. Thus there is only one tube that gets a letter, it can have any of the $2+3+6=11$ letters, as we expect from (6), so the coefficient of $x^{2}$ in $\gamma(x)$ is $11$.
@@
Next consider the coefficient of $x^{3}$ in $\gamma(x)$. There are two rooted trees on three vertices, each of which has two tubings as illustrated on the right in Figure 1. The number of upper tubes rooted at each vertex is indicated in the figure. In each case can count the number of ways to put letters at the upper tubes around each vertex via the formula of (6), but to keep this example explicit let us count directly. For the third tubing from the right, there are $11$ ways to label each upper tube since they are rooted at different vertices so there is no constraint on which letters can be chosen. For each of the other three tubings we must label a pair of tubes with the same root. If the outer of the two tubes is labelled by a letter from $\Omega_{3}$ then there is no constraint on the label of the inner of the two tubes. If the outer of the two tubes is labelled by a letter from $\Omega_{2}$ then the label of the inner of the two tubes must be from $\Omega_{2}\cup\Omega_{3}$ while if the label of the outer of the two tubes is from $\Omega_{1}$ then the label of the outer of the two tubes must also be from $\Omega_{1}$. Consequently there are $6\cdot 11+3\cdot(2+3)+2\cdot 2=85$ compatible labelings of such a tubing.
Altogether, this gives that the coefficient of $x^{3}$ in $\gamma(x)$ is $11^{2}+3\cdot 85=376$, agreeing with the explicit calculations of [11, 9].
@@
To build this system of differential equations, first consider how to recursively build tubings. As discussed in [3], by the definition of tubing, the outermost tube of any tubing is partitioned into subtubes and each of those subtubes along with all the tubes within determines a tubing of a connected subgraph of the original tree. Furthermore, take any two tubings $\tau_{1}$ and $\tau_{2}$ of plane rooted trees $t_{1}$ and $t_{2}$ respectively and then choose an insertion place in $t_{2}$. Adding an edge at the insertion point connecting to the root of $t_{1}$ builds a new tree $t$ and taking all the tubes of $\tau_{1}$ and $\tau_{2}$ along with the outer tube for $t$ gives a tubing $\tau$ of $t$, and every tubing of every rooted tree can be formed uniquely in this way. Furthermore, if a plane rooted tree has $n$ vertices then it had 2n12n-1 insertion places. Therefore, if $N(n)$ is the number of tubings of rooted trees on $n$ vertices then
@@
The differential operator $2x\frac{d}{dx}-1$ serves to pull out the coefficient 2i12i-1, and hence serves to count the insertion places. This is a special case of the general interpretation of $x\frac{d}{dx}$ as a pointing operator in enumerative combinatorics [17, Ch. A.I]. The recurrence (7) for tubings of plane rooted trees was used in [3] to give a bijection between tubings of plane rooted trees and rooted connected chord diagrams, which are classically known to have the same recurrence [20].
We will use this same construction but keep track of compatible labelings to prove the following theorem.
@@
Let $\Omega_{1},\ldots\Omega_{m}$ be disjoint alphabets with $|\Omega_{i}|=a_{i}$. Let $L_{j}(x)$ be the generating series for $\Omega_{1},\ldots\Omega_{m}$-compatible labelings of tubings of plane rooted trees for which the label on the outermost tube comes from $\Omega_{j}$. Then the anomalous dimension $\gamma(x)$ of the perturbative solution to (1) with Mellin transform (5) satisfies the following system of differential equations.
@@
As before let $\tau_{1}$ and $\tau_{2}$ be compatibly labelled tubings of plane rooted trees $t_{1}$ and $t_{2}$ respectively and choose an insertion place in $t_{2}$. Construct $\tau$ and $t$ as above by inserting $t_{1}$ into $t_{2}$ at the chosen insertion place. The labelings of the tubes of $\tau_{1}$ and $\tau_{2}$ remain compatible in $\tau$. The only thing missing is a label for the upper tube of the partition of the outermost tube, which had previously been a lower tube and is now an upper tube. Call this label the outermost label of $\tau$. The outermost label of $\tau$ is constrained to be from an alphabet $\Omega_{j}$ with $j\geq i$ where the outermost label of $\tau_{2}$ is from $\Omega_{i}$, or if $t_{2}$ is a single vertex and so has no labelled tubes then the outermost label of $\tau$ is unconstrained. Furthermore, this is the only constraint on the outermost label of $\tau$ and every compatible labeling can be obtained uniquely in this way. The labels from $\tau_{1}$ put no constraint on the outermost label of $\tau$ since they do not share roots.
As above we can encode this decomposition by generating series. Every compatible labeling of a tubing of a plane rooted tree has its outermost label from exactly one of the alphabets except for the tree with one vertex and so
@@
For a compatible labeling with the outermost label from $\Omega_{j}$, the decomposition described above tells us we have $a_{j}$ choices for the outermost label, we can take any compatible labeling of a tubing of a plane rooted tree for $\tau_{1}$, hence giving a factor of $\gamma(x)$ in the generating series, and we can take either a compatible labeling of a tree with one vertex or any compatible labeling with outermost label from $\Omega_{1}\cup\cdots\cup\Omega_{j}$ for $\tau_{2}$, where as before the differential operator $2x\frac{d}{dx}-1$ incorporates the count of the number of insertion places, and so
@@
Remark 3.6.
Very similar systems of equations were already known in the work of Marc Bellon, see equations 26 and 27 of [4]. There are minor differences, with the different equations of the system in Bellon’s case corresponding to different variables in the Mellin transform, one pole per variable, while we get the different equations in our systems from different poles in the same variable, leading to our triangular form. However, the two contexts agree on their common intersection, and so, despite being developed independently, Bellon’s work is an important predecessor of ours, and one which already appreciated the suitability of such systems for investigating asymptotics.
The main novelty of our approach is not so much the equations of Theorem 3.5 themselves, but the fact that the combinatorial interpretation means that the equations fall out automatically from the combinatorial structure – they are essentially a trivial rewriting of the recursive structure of tubings – and that each piece of the equation has a combinatorial meaning, providing insight into both specializations and generalizations.
@@
Let $\gamma_{j}:=x+\sum_{k=1}^{j}L_{k}$. For $j=1$, we have $x=\gamma_{1}-L_{1}$, and since
$L_{1}=a_{1}\gamma\left(2x\frac{d}{dx}-1\right)\gamma_{1},$
we have $x=\left[1-a_{1}\gamma\left(2x\frac{d}{dx}-1\right)\right]\gamma_{1}$. Similarly, $\gamma_{j+1}-L_{j+1}=\gamma_{j}$ implies
$\left[1-a_{j+1}\gamma\left(2x\frac{d}{dx}-1\right)\right]\gamma_{j+1}=\gamma_{j}.$
The result follows by induction, noting that $\gamma_{m}=\gamma$, by definition. ∎
@@
An advantage of the differential equation system form in Theorem 3.5 is that it provides a convenient way to compute the coefficients of the power series $\gamma(x)=\sum_{n\geq 1}g_{n}x^{n}$. To do so, we substitute power series ansätze for $\gamma(x)$ and $L_{j}(x)$ into the differential equation system. This yields a system of nonlinear difference equations for the coefficients that can be solved recursively (see Section 3.1.2 for a simple example of this procedure). We implemented this recursion as the python program gamma.py which is included in the ancillary files to the arXiv version of this article. In a short time, this program computes the numbers $g_{n}$ for various input parameters $a_{1},\ldots,a_{m}\in{\mathbb{Z}}$ up to $n\approx 1000$.
@@
Its unique power series solution (see [21, A000699]) starts with
@@
This is one of the rare cases where the prefactor $\frac{1}{e}$ is known exactly thanks to a functional equation that the associated generating function fulfills (see [6, § 7.1]).
@@
This system is equivalent to a third-order nonlinear ODE with quartic nonlinearity as in Corollary 3.7:
@@
The first coefficients of the unique power series solution (see [21, A051862]) are
@@
It is equivalent to a second-order ODE with cubic nonlinearity,